# Identifying Halos in Cosmological Simulations with Continuous Wavelet Analysis: The 2D Case

### v2

@@ Abstract
Continuous wavelet analysis is gaining popularity in science and engineering for its ability to analyze data across spatial and scale domain simultaneously. In this study, we introduce a wavelet-based method to identify halos and assess its feasibility in two-dimensional (2D) scenarios. We begin with the generation of four pseudo-2D datasets from the SIMBA dark matter simulation by compressing thin slices of three-dimensional (3D) data into 2D. We then calculate the continuous wavelet transform (CWT) directly from the particle distributions, identify local maxima that represent actual halos, and segment the CWT to delineate halo boundaries. A comparison with the traditional Friends-of-Friends (FOF) method shows that our CWT-identified halos, while containing slightly fewer particles, have smoother boundaries and are more compact in dense regions. In contrast, the CWT method can link particles over greater distances to form halos in sparse regions due to its spatial segmentation scheme. The spatial distribution and halo power spectrum of both CWT and FOF halos demonstrate substantial consistency, validating the 2D applicability of CWT for halo detection. Our identification scheme operates with a linear time complexity of $\mathcal{O}(N)$, suggesting its suitability for analyzing significantly larger datasets in the future.
software: NumPy (https://numpy.org/), SciPy (https://scipy.org/), Matplotlib (https://matplotlib.org/), Jupyter Notebook (https://jupyter.org/)
@@ 1 Introduction
In modern cosmology, it is assumed that our universe was highly homogeneous in its early stages. Over time, these initial minute density perturbations evolve into a variety of structures, such as dark matter halos, galaxies, clusters, and voids. This evolution occurs through linear and nonlinear processes, driven by long-range gravity (Mo et al., 2010). To study the universe’s evolution, observations of astronomical objects at high redshifts and extremely large distances are necessary. These observations are complemented by numerical simulations, which are indispensable tools for cosmological research. Such simulations begin with an initial condition and then solve the $N$-body dynamic equations or hydrodynamic equations to simulate the universe’s evolution. Within these simulations, the universe’s density is represented by particles, typically dark matter and baryonic matter particles, which perform a Poisson sampling of the simulated density field (Peebles, 1980).
@@
With the rapid advancement of computer science, cosmological simulations have progressed from simple $N$-body models involving pure dark matter bound by gravity (Press & Schechter, 1974) to sophisticated adaptive moving-mesh magneto-hydrodynamic simulations (Nelson et al., 2019) and meshless finite-mass hydrodynamic simulations (Davé et al., 2019), which incorporate processes of feedback and interactions. These advanced simulations enable us to model the evolution of the universe from redshifts of a few hundred to the present day within a simulation box spanning millions or even billions of parsecs in comoving units. While the simulations align with observations in various aspects - such as the existence of cosmic web, the characteristics of stars, and the gas content of galaxies - the analysis of simulated particle data remains challenging. For instance, when determining statistical properties like the power spectrum, particles are commonly assigned to a regular grid using window functions, such as the Count-in-Cell (CIC) or Triangular-Shaped-Cloud (TSC) methods (see Hockney & Eastwood, 1981). Moreover, constructing a halo catalog to study the clustering of matter in the simulated universe necessitates identifying individual halos at the particle data limit and considering a cluster of particles as a single halo.
@@
For this purpose, we can employ various halo finders, such as Amiga Halo Finder (AHF) (Knollmann & Knebe, 2009), adaptive spherical overdensity halo finder (ASOHF) (Planelles & Quilis, 2010) and ROCKSTAR halo finder (Behroozi et al., 2013), to group particles into a halo. Among them, the most widely used algorithm is Friends-of-Friends (FOF) method (Davis et al., 1985). The FOF method calculates a linking length based on the particle number density and then connects particles whose distance is less than this length. Despite its simplicity, this method is popular for identifying dark halos from particle data due to its ability to provide an accurate measure of halo distribution. However, the FOF method has limitations, in that it has a time complexity of $\mathcal{O}(N\log N)$ when utilizing a minimal spanning tree for organizing particle data, and it employs an artificial parameter $b$ in the linking length. Worst of all, it does not offer a clear boundary definition, leading to the ‘linking bridges’ problem, where two distinct halos may be erroneously linked if a chain of closely spaced (less than the linking length) particles exists between them (Springel et al., 2001).
@@
In addition to the FOF method, an alternative for analyzing clustering is the wavelet transform method, first introduced by Grossmann et al. (1985). This technique is extensively utilized in astronomy and cosmology to identify structures, such as analyzing clusters in X-ray data (e.g. Vavilova & Babyk, 2018; Freeman et al., 2002), infrared data (Wang et al., 2008), and gamma-ray data (Ciprini et al., 2007). It is also used to study how discreteness effects in $N$-body simulations affect the statistical diagnostics of the cosmological density field (Romeo et al., 2008), to trace the time evolution of the cosmic web (Einasto et al., 2011), or to identify sunspots (Djafer et al., 2012), and to characterize coronal mass ejections (González-Gómez et al., 2010). Furthermore, the wavelet method is employed to detect Baryon Acoustic Oscillation (BAO) structures (Arnalte-Mur et al., 2012), ridges of spiral arms (Patrikeev et al., 2006), and point source contamination in Cosmic Microwave Background (CMB) observations (Cayón et al., 2000).
@@
There are two distinct wavelet methods, the discrete wavelet transform (DWT) and the continuous wavelet transform (CWT). The DWT decomposes a signal into a set of orthogonal wavelet bases, offering a compact representation. In contrast, the CWT involves the convolution of the signal with a wavelet that is scaled by a mother wavelet. The CWT computation is typically performed on a grid, or image pixels in observations, and uses the Fast Fourier Transform (FFT), exploiting the property that convolution in real space is equivalent to multiplication in Fourier space (e.g. Arnalte-Mur et al., 2012; Grebenev et al., 1995; Barnard et al., 2004; Kazakevich et al., 2004).
@@
In structure identification, the DWT is commonly applied using the ‘$\grave{{\rm a}}$ trous’ algorithm, which filters an image with scaling functions at doubled magnification scales and then decomposes it into wavelets at various scales, along with a residual (e.g. Bijaoui & Giudicelli, 1990; Pagliaro et al., 1999; Ellien et al., 2021; Lazzati et al., 1999; Moretti et al., 2004). This algorithm, distinct from the standard DWT, employs a non-orthogonal wavelet basis, enhancing its ability to localize structures (Chereul et al., 1998). It enables rapid computation of grid DWTs at multiple scales with a constant resolution but fixes the scaling to a doubly magnified set of scales.
@@
Compared with the DWT, the CWT typically demands more time and CPU resources for any given data. However, for the discrete particle data encountered in simulations, an alternative method exists to compute the CWT with a time complexity of $\mathcal{O}(N)$. This approach views the density field as a sum of Dirac delta functions, which means that the corresponding CWT is the convolution of these Dirac functions with a wavelet, yielding a sum of wavelets. Although some authors describe this method as a DWT (e.g. Bendjoya et al., 1991; Slezak et al., 1993; Flin & Krywult, 2006), the outcome of this convolution is inherently continuous, as its value can be determined at any position and scale, irrespective of resolution.
@@
In this paper, we introduce a novel wavelet-based method for two-dimensional (2D) halo identification, visualize the results, and compare them with the traditional FOF method. The paper is organized as follows. In Section 2, we provide a brief overview of the CWT method, the FOF method, the SIMBA data utilized, and the derived pseudo-2D data. Section 3 details our multi-scale CWT halo identification approach. In Section 4, we calculate the statistical significance and investigate the impact of resolution on our method. Section 5 presents our identification findings and compares them with the FOF halo identifications. Finally, we offer conclusions and discuss the implications of our results in Section 6.
@@ 2.1 The Continuous Wavelet Transform
where $\Psi(w,\mathbf{x})$ is a scaled version of the so-called mother wavelet $\Psi(\mathbf{x})$, as follows
@@
We introduce the scale parameter $w$ instead of the traditional scale indicator $a$, with the relationship between them being $w=1/a$ (Wang et al., 2022; Wang & He, 2022). For discrete particle data from simulations, their density field can be expressed as the sum of the Dirac delta function $\delta^{(\rm D)}(\mathbf{x})$ located at their positions, weighted by their mass,
@@
where $m_{i}$, $\mathbf{x}_{i}$ represent the mass and position vector of the $i$-th particle, respectively. Thus, the CWT of the particle density field is the convolution of the Dirac function with our wavelet, which is simply the translated wavelet itself
@@
It is almost impossible to analyse Equation (2.1) analytically. In this context, we must analyse millions of functions simultaneously (the number of functions equals the number of particles). Instead, we discretize Equation (2.1) onto a regular grid, that is, we calculate the value of Equation (2.1) at specific grid points $\mathbf{x}_{\rm g}$ to represent it:
@@
Even if we employ isotropic wavelets, they can still be used to detect anisotropic structures (Vavilova & Babyk, 2018). To simplify our calculations, we opt for isotropic wavelets with compact support, or at least with an approximate compact support. Furthermore, if a wavelet is not orthogonal, a particular structure might exhibit local maxima at several scales, rather than solely at the scale that most closely matches the structure’s size. This could lead to ambiguity when identifying halos within a multi-scale analysis (Hayn et al., 2012; Ellien et al., 2021). However, it is not feasible for a wavelet to simultaneously possess compact support, isotropy, and orthogonality (Daubechies, 1992). Therefore, we select the MH wavelet for its isotropy and approximate compact support (Wang & He, 2021). The CWT has a notable property: the CWT of any constant or linear component of a signal is zero. This characteristic provides us with a method to distinguish between different structures, even in highly dense environments, and offers a natural delineation.
Figure 1: The density field of our pseudo-2D data in the CIC grid, from top left to bottom right are slices at ${\rm z}=11.4$, $28.4$, $38.0$ and $45.7\,{h^{-1}{\rm Mpc}}$, respectively. The thickness of all the four slices is $\Delta{\rm z}=0.1\,{h^{-1}{\rm Mpc}}$.
Figure 2: The same halo is depicted with directly assigned particles (left) and weighted assigned particles (right). The halo assigned directly exhibits a very sharp boundary, whereas the weighted assignment results in a much smoother boundary.
@@ 2.2 Friends-of-Friends Algorithm
The FOF method groups particles into different halos (Davis et al., 1985). The grouping process can be summarized as follows: First, an arbitrary particle is selected. Then search for its nearby particles and links any particle whose distance to the selected one is less than the linking length, $l_{\rm link}$; these are termed the ‘friends’ of the initial particle. Subsequently, the search continues for the ‘friends’ of each newly linked particle. This linking continues to include additional ‘friends’ whose distance to any current ‘friend’ is also less than $l_{\rm link}$. The process repeats, finding ‘friends of friends’, until no new particles can be linked under this criterion. These interconnected particles ultimately form a FOF group.
After that, select another arbitrary particle that is not already in any group and search for its ’friends’ to form a new group. Repeat this process until all particles are assigned to groups, even if some groups contain only one particle. Finally, exclude groups that have fewer than a specified threshold number of particles, $N_{\rm th}$; for example, $N_{\rm th}=5$. This results in the classification of only those halos that contain five or more particles.
@@
where $V$ is the volume of the simulation box, $N$ is the number of particles, and $b=0.2$ so that the average density in the halo is 180 times the average density in the box (More et al., 2011). In the 2D case, Equation (9) can be rewritten as
@@
where $S$ is the area of the simulation box and we still set $b=0.2$. However, the choice for the latter, $N_{\rm th}$, is much more arbitrary. Although $N_{\rm th}$ means the mass of the smallest halo, there is no consensual definition [^1] Someone authors set $N_{\rm th}=32$ (see http://swift.dur.ac.uk/docs/FriendsOfFriends/algorithm_description.html), while others have different definitions (e.g. ytastro toolkit has the default $N_{\rm th}=5$) (Turk et al., 2011)..

### v3

@@ Abstract
Continuous wavelet analysis is gaining popularity in science and engineering for its ability to analyze data across both spatial and scale domains simultaneously. In this study, we introduce a wavelet-based method for halo identification and assess its feasibility in two-dimensional (2D) scenarios. We begin by generating four pseudo-2D datasets from the SIMBA dark matter simulation by compressing thin slices of three-dimensional (3D) data into 2D. Subsequently, we calculate the continuous wavelet transform (CWT) directly from the particle distributions, identify local maxima that represent actual halos, and segment the CWT to delineate halo boundaries. A comparison with the traditional friends-of-friends (FOF) method shows that while our CWT-identified halos contain slightly fewer particles, they have smoother boundaries and are more compact in dense regions. In contrast, the CWT method can link particles over greater distances to form halos in sparse regions due to its spatial segmentation scheme. The spatial distribution and halo power spectrum of both CWT and FOF halos demonstrate substantial consistency, validating the 2D applicability of CWT for halo detection. While our identification scheme has been tested solely in a limited 2D context and has shown some performance limitations, its linear time complexity of $\mathcal{O}(N)$ and consistency with the FOF method suggest its suitability for analyzing significantly larger datasets in the future and its potential to be extended to the 3D case.
software: NumPy (van der Walt et al., 2011; Harris et al., 2020) [^3] https://numpy.org/, SciPy (Virtanen et al., 2020) [^4] https://scipy.org/, Matplotlib (Hunter, 2007) [^5] https://matplotlib.org/, Jupyter Notebook [^6] https://jupyter.org/, yt project (Turk et al., 2011) [^7] https://github.com/yt-project
@@ 1 Introduction
In modern cosmology, it is assumed that our Universe was highly homogeneous in its early stages. Over time, these initial minute density perturbations evolved into a variety of structures, such as dark matter halos, galaxies, clusters, and voids. This evolution occured through both linear and nonlinear processes, driven by long-range gravity (Mo et al., 2010). To study the Universe’s evolution, it is essential to conduct observations of astronomical objects at high redshifts and over vast cosmic distances. These observations are complemented by numerical simulations, recognized as indispensable tools for cosmological research. Such simulations begin with initial conditions and then solve the $N$-body dynamic equations or hydrodynamic equations, thereby modeling the Universe’s evolution. Within these simulations, the density of the Universe is usually represented by particles, typically dark matter and baryonic matter particles (in the case of Lagrangian codes) (Vogelsberger et al., 2020), which perform a Poisson sampling of the simulated density field (Peebles, 1980).
@@
With the rapid advancement of computer science, cosmological simulations have progressed from rudimentary $N$-body models involving pure dark matter bound by gravity (Press & Schechter, 1974) to sophisticated adaptive moving mesh magnetohydrodynamic simulations (Nelson et al., 2019) and meshless finite mass hydrodynamic simulations (Davé et al., 2019), which incorporate the processes of feedback and interactions. These advanced simulations enable us to model the evolution of the Universe from redshifts of a few hundred to the present day, within a simulation box spanning millions or even billions of parsecs in comoving units. While the simulations align with observations in various aspects - such as the existence of the cosmic web, the characteristics of stars, and the gas content of galaxies - the analysis of simulated particle data remains challenging. For instance, when determining statistical properties like the power spectrum, particles are commonly assigned to a regular grid using window functions, such as the cloud-in-cell (CIC) or triangular-Shaped cloud methods (see Hockney & Eastwood, 1981). Moreover, constructing a halo catalog to study the clustering of matter in the simulated universe necessitates identifying individual halos at the particle data level and considering a cluster of particles as a single halo.
For this purpose, we can employ various halo finders, such as the Amiga Halo Finder (Knollmann & Knebe, 2009), the adaptive spherical overdensity halo finder (Planelles & Quilis, 2010; Vallés-Pérez et al., 2022), HOP (Eisenstein & Hut, 1998), and ROCKSTAR (Behroozi et al., 2013). Halo finders can generally be categorized into two types: density peak locators and particle collectors. The former, based on the spherical-overdensity method (Press & Schechter, 1974), identifies the density peaks and subsequently selects particles encircling these peaks to constitute halos. The latter, exemplified by the friends-of-friends (FOF) method (Davis et al., 1985), links particles that satisfy specific conditions together (e.g., the distance below the linking length in FOF, or toward the densest direction in HOP).
Generally speaking, a halo finder typically involves the following processes: identifying halo candidates and excluding spurious halos, collecting particles associated with these candidates, determining the halo center, conducting a self-boundness check and unbinding free particles, delineating the halo boundary, and tracing halos across different snapshots. Due to the extreme variability among different halo identification methods, not all halo finders undertake each of the aforementioned processes. For instance, some FOF-based halo finders do not generate a halo candidate catalog. The standard FOF halo finder excludes the unbinding process, and most early halo finders do not track the time evolution of halos. Despite the existence of these disparate halo finders, a degree of uncertainty persists. A primary source of uncertainty lies in defining the halo boundary, which is closely related to the fundamental question of what constitutes a halo. The criteria for this definition vary significantly across different halo finders (Knebe et al., 2011, 2013).
@@
Among them, the most widely used algorithm is the FOF method, which calculates a linking length based on the particle number density and then connects particles that are closer than this length. Despite its simplicity, this method is popular for identifying dark halos from particle data, as it offers a reliable measure of halo distribution. However, the FOF method has limitations in that it has a time complexity of $\mathcal{O}(N\log N)$ when utilizing a minimal spanning tree for organizing particle data, and it employs an artificial parameter $b$ in the linking length. Worst of all, it lacks a clear boundary definition, leading to the ‘linking bridges’ problem, erroneously connecting two distinct halos if there is a chain of closely spaced particles (below the linking length) between them (Springel et al., 2001).
@@
In addition to the FOF method, the wavelet transform method, first introduced by Grossmann et al. (1985), offers an alternative approach to clustering analysis. This technique is extensively utilized in astronomy and cosmology to identify structures, such as analyzing clusters in X-ray data (e.g., Vavilova & Babyk, 2018; Freeman et al., 2002), infrared data (Wang et al., 2008), and gamma-ray data (Ciprini et al., 2007). Additionally, it serves to investigate how discreteness effects in $N$-body simulations affect the statistical diagnostics of the cosmological density field (Romeo et al., 2008), trace the time evolution of the cosmic web (Einasto et al., 2011), identify sunspots (Djafer et al., 2012), and characterize coronal mass ejections (González-Gómez et al., 2010). Furthermore, the wavelet method is employed to detect baryon acoustic oscillation structures (Arnalte-Mur et al., 2012), locate ridges of spiral arms (Patrikeev et al., 2006), and identify point source contamination in cosmic microwave background observations (Cayón et al., 2000).
@@
There are two distinct wavelet methods: the discrete wavelet transform (DWT) and the continuous wavelet transform (CWT). The DWT decomposes a signal into a set of orthogonal wavelet bases, offering a compact representation. In contrast, the CWT involves the convolution of the signal with a wavelet that is scaled by a mother wavelet. The CWT computation is typically performed on a grid, or on image pixels in observations. It employs the fast Fourier transform and exploits the property that convolution in real space is equivalent to multiplication in Fourier space (e.g., Arnalte-Mur et al., 2012; Grebenev et al., 1995; Barnard et al., 2004; Kazakevich et al., 2004).
@@
In structure identification, the DWT is frequently implemented via the ‘$\grave{{\rm a}}$ trous’ algorithm. This algorithm filters an image with scaling functions at doubled magnification scales and then decomposes it into wavelets at various scales, along with a residual (e.g., Bijaoui & Giudicelli, 1990; Pagliaro et al., 1999; Ellien et al., 2021; Lazzati et al., 1999; Moretti et al., 2004). This algorithm, distinct from the standard DWT, employs a nonorthogonal wavelet basis, enhancing its ability for structural localization (Chereul et al., 1998). It enables rapid computation of grid DWTs at multiple scales with a constant resolution, but the scaling is fixed to a doubly magnified set of scales.
@@
Compared with the DWT, the CWT typically demands more time and CPU resources. However, when dealing with discrete particle data from simulations, an alternative method exists to compute the CWT with a time complexity of $\mathcal{O}(N)$. This approach views the density field as a sum of Dirac delta functions, meaning that the corresponding CWT is the convolution of these Dirac functions with a wavelet, yielding a sum of wavelets. Although some authors describe this method as a DWT (e.g., Bendjoya et al., 1991; Slezak et al., 1993; Flin & Krywult, 2006), the outcome of this convolution is inherently continuous, as its value can be determined at any position and scale, regardless of resolution.
@@
In this paper, we introduce a novel wavelet-based method for two-dimensional (2D) halo identification, visualize the results, and compare them with the traditional FOF method. The paper is organized as follows. In Section 2, we provide a brief overview of the CWT method, the FOF method, the utilized SIMBA dataset, and the derived pseudo-2D dataset. Section 3 details our multiscale CWT halo identification approach. In Section 4, we calculate the statistical significance and investigate the impact of resolution on our method. Section 5 presents our identification findings and compares them with the FOF halo identifications. Finally, we offer conclusions and discuss the implications of our results in Section 6. The actual performance of our current program is presented in the Appendix A.
@@ 2.1 The CWT
where $\Psi(w,\mathbf{x})$ is a scaled version of the so-called mother wavelet $\Psi(\mathbf{x})$, as follows:
@@
We introduce the scale parameter $w$ instead of the traditional scale indicator $a$, with the relationship between them being $w=1/a$ (Wang et al., 2022; Wang & He, 2022). For discrete particle data from simulations, their density field can be expressed as the sum of the Dirac delta functions $\delta^{(\rm D)}(\mathbf{x})$ located at their positions and weighted by their masses:
@@
where $m_{i}$, $\mathbf{x}_{i}$ represent the mass and position vector of the $i$ th particle, respectively. Thus, the CWT of the particle density field is the convolution of the Dirac function with our wavelet, which is simply the translated wavelet itself
@@
It is almost impossible to analyze Equation (2.1) analytically. In this context, we must analyze millions of functions simultaneously (the number of functions equals the number of particles). Instead, we discretize Equation (2.1) onto a regular grid; that is, we calculate the value of Equation (2.1) at specific grid points $\mathbf{x}_{\rm g}$ to represent it:
@@
Even when employing isotropic wavelets, they can still be used to detect anisotropic structures (Vavilova & Babyk, 2018). To simplify our calculations, we opt for isotropic wavelets with compact support, or at least with an approximate compact support. Furthermore, if a wavelet is not orthogonal, a particular structure might exhibit local maxima at several scales, rather than solely at the scale that most closely matches the structure’s size. This could lead to ambiguity when identifying halos within a multiscale analysis (Hayn et al., 2012; Ellien et al., 2021). However, it is not feasible for a wavelet to simultaneously possess compact support, isotropy, and orthogonality (Daubechies, 1992). Therefore, we select the MH wavelet for its isotropy and approximate compact support (Wang & He, 2021). The CWT has a notable property: the CWT of any constant or linear component of a signal is zero. This characteristic provides us with a method to distinguish between different structures, even in highly dense environments, and offers a natural delineation.
Figure 1: The density field of our pseudo-2D data in the CIC grid. From top left to bottom right are slices at ${\rm z}=11.4$, $28.4$, $38.0$ and $45.7\,{h^{-1}{\rm Mpc}}$, respectively. The thickness of all the four slices is $\Delta{\rm z}=0.1\,{h^{-1}{\rm Mpc}}$ and all the grids are made of $512^{2}$ cells.
Table 1: List of the parameters used in This paper, along with Their Meanings and Definitions.
Parameter Meaning Definition
a The physical scale of the wavelet None
w The scale parameter $1/a$
$k_{w}$ The dimensionless scale parameter $\frac{L_{\rm box}}{1000}w$
$N_{\rm g}$ The resolution of the grid ${\rm Int}(1000\times k_{w})+500$
$N_{\rm W}$ The resolution of the wavelet ${\rm Int}(2\times(10/k_{w}+10))+1$
$N_{\rm cs}$ The resolution of cross-scale comparison 1000
$N_{\rm th}$ The least particle number in halo 5
@@ 2.2 Friends-of-Friends Algorithm
The FOF method categorizes particles into distinct halos (Davis et al., 1985). The grouping process can be summarized as follows: First, an arbitrary particle is selected. Then, this algorithm searches for its nearby particles and links any particle whose distance to the selected one is less than the linking length, $l_{\rm link}$; these are termed the friends of the initial particle. Subsequently, the search continues for the friends of each newly linked particle. This linking continues to include additional friends whose distance to any current friend is also less than $l_{\rm link}$. The process iterates to identify friends of friends until no further particles meet the linking criterion. These interconnected particles ultimately form a FOF group.
After that, select another arbitrary particle that is not already in any group and search for its friends to form a new group. Repeat this process until all particles are assigned to groups, even if some groups contain only one particle. Finally, exclude groups that have fewer than a specified threshold number of particles, $N_{\rm th}$; for example, $N_{\rm th}=5$. This results in the classification of halos comprising five or more particles.
@@
where $V$ denotes the volume of the simulation box, $N$ is the number of particles, and $b=0.2$ so that the average density in the halo is 180 times the average density in the box (More et al., 2011). In the 2D case, Equation (9) can be rewritten as
@@
where $S$ is the area of the simulation box and we still set $b=0.2$. However, the choice for the latter, $N_{\rm th}$, is much more arbitrary. Although $N_{\rm th}$ represents the mass of the smallest halo, there is no consensual definition [^1] Some authors set $N_{\rm th}=32$ (see http://swift.dur.ac.uk/docs/FriendsOfFriends/algorithm_description.html), while others have different definitions (e.g., ytastro toolkit has the default $N_{\rm th}=5$) (Turk et al., 2011)..